five times the measure of an angle is 48 more than the measure of its supplement. find the measure if the angle and its supplement.
I started my equation as 5x=180+x+48 then i combined like terms and got 6x=180=48 and for x i got 38. Then i was lost

Answer by mananth(16946)

angle = x
supplement = (180-x)
5x=(180-x)+48
5x=180-x+48
5x+x=180-48
6x=228
/6
x=38 deg the angle
supplement = 180-x
180-38
142 deg the supplement
m.ananth@hotmail.ca